OpenStudy (anonymous):

If water flows at 20ft^3/sec and there are 7.5 gallon is a cubic foot, how many minutes will it take to fill a 75,000 gallon tank?

hero (hero):

7.5g / xg = 1ft^3/20ft^3 x = 150 150g/75000g = 20ft^3/yft^3 Solve for y

OpenStudy (anonymous):

im in a timed quiz

OpenStudy (anonymous):

i dont have the time to slove for y can u help me out pls

OpenStudy (anonymous):

im working on other problems

hero (hero):

Well, I get y = 100, but it doesn't seem right because that would mean 100 seconds and converted to minutes would be 1 min 40 secs so something is off I think.

OpenStudy (anonymous):

my options are 4, 5000, 50, 8.5

hero (hero):

Okay, let me try again, I just notices something I overlooked before

OpenStudy (anonymous):

i have like 5 mins like can you please help me

hero (hero):

8.3 min

hero (hero):

Was it 8.3 or 8.5?

OpenStudy (anonymous):

thanks you so much. you are definately a hero

OpenStudy (anonymous):

If there are four quarts in a gallon and six gallons in a tank and I drive 20 miles on a gallon, how far can I drive on 12 quarts? can you look at that quickly i think i know the answer just want to make sure

OpenStudy (anonymous):

hahaha i posted the same question sorry

OpenStudy (anonymous):

If there are four quarts in a gallon and six gallons in a tank and I drive 20 miles on a gallon, how far can I drive on 12 quarts?

OpenStudy (anonymous):

never mind this is the right question can you look at it quickly

hero (hero):

Jinnie, say something...

hero (hero):

Thanks for leaving me hanging

hero (hero):

I think it was 60 miles instead of 50. Sorry, I was rushing to try to help you.